The demand and supply curves for the market are given by: Demand: Qd = 16000 – 24P Supply: Qs = 2000 + 32P

A maximum price of $200 is proposed.

(ii) Calculate the equilibrium quantity that would prevail without the maximum price
Answer #1

As per the question

Market demand function Qd = 16000 -24P

Market supply function Qs = 2000 + 32P

Proposed maximum price = $200

Equilibrium quantity without the maximum price = ?

At the market equilibrium the Qd = Qs

16000 -24P = 2000 + 32P

16000 – 2000 = 32P + 24P

14000 = 56P

56P=14000

P=14000/56 =250

Substituting the value of P=250, in market demand function to the value of equilibrium quantity (Q)

Qd = 16000 -24(250) = 16000 – 6000 = 10000

So without the maximum price, the equilibrium quantity that would prevail in market is 10000 units